# WiFo-M2: Plug-and-Play Multi-Modal Sensing via Foundation Model to Empower Wireless Communications

### v1

@@ V-C1 Beam Prediction
  • $\bullet$
    Experiment Settings: The training hyperparameters for Vision-BP and MM-BP schemes are kept identical to those used in [7] and [29]. The proposed WiFo-M2-based BP scheme is trained with a batch size of $32$ and an initial learning rate of $2\times 10^{-4}$, without any learning rate decay.
  • $\bullet$
    WiFo-M2’s Performance: We evaluate the performance of BP task using the top-$k$ accuracy metric. This metric measures the percentage of test samples for which the ground-truth optimal beam is included in the model’s $k$ highest-ranked predictions. In our experiments, we report results for $k=1,3,5$ to comprehensively assess the prediction capability.
    As shown in Fig. 6, across all $7$ datasets, the proposed WiFo-M2 models exhibit highly competitive prediction accuracy even though the backbones are completely frozen. The WiFo-M2-Img scheme attains noticeably higher Top-1 prediction accuracy than Vision-BP scheme on L4, L5 and L6, whereas on the remaining four datasets its performance is only slightly lower than that of Vision-BP. This demonstrates that the WiFo-M2-Img backbone acquires channel-aware knowledge extracted during pre-training. The effect is more pronounced on LiDAR-based BP task: WiFo-M2-LiDAR substantially surpasses the fully-trained MM-BP-LiDAR baseline on every dataset, which indicates that the pre-trained WiFo-M2-LiDAR captures spatial patterns that are highly predictive of the optimal beam. When both modalities are available, WiFo-M2 delivers virtually the same performance as the MM-BP scheme, with a mean Top-1 gap of only 2%. In summary, the BP results confirm that the ContraSoM pre-training aligns visual and LiDAR features with the wireless channel domain. Consequently, high-quality beam selection can be achieved with minimal retraining overhead.
V-C2 Channel Estimation
  • $\bullet$
    Problem Description: CE task refers to the process in which the BS estimates the channel $\mathbf{h}$ at the pilot-placed subcarriers. We assume that the BS adopts the hybrid (analog/digital) structure and is equipped with $N_{\rm{RF}}$ RF chains. The digital and analog precoding matrix are represented by $\mathbf{F}_{\rm{B}}\in\mathbb{C}^{N_{\rm{RF}}\times N_{\rm{RF}}}$ and $\mathbf{F}_{\rm{R}}\in\mathbb{C}^{N\times N_{\rm{RF}}}$ with $\|\mathbf{F}_{\rm{R}}\mathbf{F}_{\rm{B}}\|_{\rm{F}}^{2}=N_{\rm{RF}}$. $\mathbf{F}_{\rm{B}}$ is set as an identity matrix $\mathbf{I}_{N_{\rm{RF}}}$ during the CE stage. The quantization bit number of the phase shifters adopted at BS is set to $B=3$. Each element of $\mathbf{F}_{\rm{R}}$ is randomly chosen from the set $\{e^{\frac{j2\pi n}{Q}},n=1,2,\dots,Q\}$, where $Q\triangleq 2^{B}=8$. During the channel estimation stage, the single-antenna user transmits known pilots $\mathbf{x}\in\mathbb{C}^{1\times T}$ over $T$ successive time slots. The received signal sequence at BS $\mathbf{Y}\in\mathbb{C}^{N_{\rm RF}\times T}$ can then be expressed as:
@@ V-C2 Channel Estimation
    where $\mathbf{n}$ denotes the additive white Gaussian noise with i.i.d. entries of zero mean and variance $\sigma^{2}$. Upon receiving the pilots, BS employs a certain CE network to estimate the CSI: $\hat{\mathbf{h}}=\mathcal{Q}_{\text{CE}}\bigl(\mathbf{Y},\mathcal{G}(\mathbf{M}))$, where $\mathcal{Q}_{\rm CE}(\cdot)$ represents a certain CE scheme and $\mathcal{G}(\cdot)$ denotes the WiFo‑M2 model.
  • $\bullet$
@@
  • $\bullet$
    WiFo-M2’s Performance: The normalized mean square error (NMSE) is used as the estimation performance metric given by ${\rm NMSE}={{\|\hat{\mathbf{H}}-\mathbf{H}\|}_{\rm{F}}^{2}}/{{\|\mathbf{H}\|}_{\rm{F}}^{2}}$. We evaluates the channel estimation performance of CENN and FCDAMP schemes under varying SNR conditions. As shown in Fig. 7, the WiFo-M2-enhanced variants consistently outperform the original CENN and FCDAMP models across most datasets and SNRs, demonstrating the general utility of the OOB channel features. Furthermore, the multi-modal fusion (WiFo-M2) typically yields the most robust and significant gains, with LiDAR-only and image-only enhancements providing comparable improvements in most scenarios. On the L3 dataset with a large $256$-antenna array and limited pilots, the performance gains are less pronounced and sometimes irregular. This aligns with the expectation that under extremely sparse measurements and high-resolution array requirements, the coarse prior provided by multi-modal sensing becomes less effective in refining the estimation.
@@
V-C3 Channel Interpolation
  • $\bullet$
    Problem Description: During the channel estimation stage, pilots are placed only on a subset of subcarriers, and a subset of BS’s antenna elements is activated [^1] In our CI task, we assume that a subset of antennas are provided with coarse channel estimates, while the CI schemes are required to recover the full CSI across all antennas. Although this setup does not reflect a practical system configuration, it serves as an effective testbed for evaluating the evaluated CI schemes’ ability to perform spatial super‑resolution reconstruction.. Upon receiving pilots, BS estimates the channel at pilot positions using the Least Squares (LS) algorithm. All the available channel estimates forms the low-resolution (LR) channel observation, denoted by $\hat{\mathbf{H}}^{\rm LR}\in\mathbb{C}^{K_{\rm p}\times N_{\rm p}}$. Specifically, if pilots are placed every $k$ subcarriers and every $x$ antenna elements, the LR observation can be expressed as:
@@ V-C3 Channel Interpolation
  • $\bullet$
    Experiment Settings: We assume that the pilot placement interval in subcarrier dimension and the sampling ratios for antenna dimensions are set to $k=K/8$ and $x=N/8$, respectively. That is, coarse channel estimates are available on only $1/8$ of the subcarriers and $1/8$ of the antennas, respectively. For network hyperparameters, the SR-CI scheme uses a learning rate of $1\times 10^{-2}$, the LPCCNet scheme uses $1\times 10^{-4}$. The batch size for both schemes is $32$.
  • $\bullet$
    WiFo-M2’s Performance: The performance gains of WiFo-M2 models over baseline LPCCNet and SR-CI schemes are illustrated in Fig. 8. Among the three variants, WiFo-M2 consistently achieves the most stable and significant improvements across different datasets. This demonstrates the effectiveness of multi-modal fusion in enhancing CI accuracy. Note that on the L3 dataset the WiFo-M2 models introduce a marginal performance degradation. A likely reason is that the L3 dataset configures a large antenna array, which translates into a very fine angular-resolution requirement for CI. In such a regime, the spatial side information from multi-modal sensing remain too coarse or imperfectly aligned with the complex channel, so the supplementary features cannot improve the high-precision estimates demanded by LPCCNet. Moreover, the performance gains brought by WiFo-M2 are especially pronounced on the L4 and L6 datasets. L4 employs a BS with a small antenna array, while L6 is configured with relatively fewer subcarriers. Under these reduced spatial- or frequency-resolution settings, the OOB features provided by WiFo-M2 are just sufficient to recover the correspondingly lower-dimensional CSI, leading to the sharp performance boost.
V-C4 Channel Prediction
  • $\bullet$
    Problem Description: In frequency division duplex (FDD) systems, where uplink-downlink reciprocity does not hold, the BS relies on uplink channel feedback to obtain downlink CSI. However, the feedback overhead scales linearly with the number of antennas and devices, making it infeasible for massive MIMO systems. Given that the uplink and downlink transmissions experience the same physical environment, it is reasonable to utilize the uplink CSI measured at the BS to predict the downlink CSI. Similarly to [37, 38], the first half of the subcarriers are considered as the uplink channel, while the remaining half are treated as the downlink channel. Let $K_{\rm in}$ be the number of subcarriers in the first half, and $K_{\rm out}$ represents the number of subcarriers in the second half. Then, the uplink channel $\mathbf{H}_{\rm d}$ and downlink channel $\mathbf{H}_{\rm u}$ can be expressed as:
@@ V-C4 Channel Prediction
    The CP task can be formulated as: $\mathbf{H}_{\rm u}=\mathcal{Q}_{\rm CP}({\mathbf{H}}_{\rm d},\mathcal{G}(\mathbf{M}))$, where $\mathcal{Q}_{\rm CP}(\cdot)$ represents a certain CP scheme.
  • $\bullet$
    Existing CP schemes: We selected two existing CP schemes, namely WiFo [21] and Transformer-based scheme [39]. WiFo [21] is a wireless FM designed for time/frequency-domain CP, which is pretrained via several self-supervised training tasks. Transformer-based scheme [39] is a parallel CP scheme using Transformer.
@@
  • $\bullet$
    WiFo-M2’s Performance: The performance enhancements provided by WiFo-M2 models on top of baseline WiFo and Transformer-based schemes are summarized in Table III. The NMSE is also used as the performance metric. The WiFo-M2 model leveraging multi-modality demonstrates its superiority by delivering highest gains for both CP baselines across most datasets, underscoring the benefit of cross-modal feature fusion. Consistent with observations in CI task, performance gains are particularly significant on datasets L4 and L6. Notably, WiFo-M2’s effectiveness is inherently bounded by the convergence of the baseline scheme itself. For instance, on dataset L5, the Transformer-based baseline fails to achieve downlink CSI prediction, and correspondingly, WiFo-M2 provides no enhancement. This indicates that the supplementary features cannot compensate for fundamental algorithmic limitations in the baseline model.
@@ V-D Ablation Study
Due to page limits, we present ablation results for several representative transceiver design modules, as shown in Table IV. First, removing the temporal feature extrapolation function in WiFo-M2 leads to a performance drop, confirming that extrapolating fine-grained future features effectively bridges the sparsity of sensory data. The NMSE results reported are averaged over the L1–L7 datasets. The ablation study results demonstrate the effectiveness of each proposed component. Second, replacing the ContraSoM pretrained weights with generic ImageNet‑pretrained weights significantly degrades performance (e.g., from 9.63-9.63 to 7.55-7.55 on LPCCNet), highlighting the advantage of the supplementary channel features over general‑purpose visual features. Third, removing the DA design from WiFo-M2-LiDAR causes a performance decline, validating that the proposed feature‑space diffusion process enhances LiDAR representation learning.
@@ V-E1 S1 Dataset
The S1 dataset provides measured image and LiDAR point cloud data, as well as the receive power with beams. Due to the absence of CSI, only the performance of the BP task is evaluated on the S1 dataset. The beam prediction task on the S1 dataset is relatively straightforward due to the 16-element antenna array. The Vision-BP baseline achieves Top-1/3/5 accuracies of 91.0%, 98.7%, and 100%, while the MM-BP scheme reaches 92.2%, 100%, and 100%. Notably, the WiFo-M2-img model, with its pre-trained backbone frozen and only a final MLP layer fine-tuned, attains 91.0%, 100%, and 100%, matching or surpassing the strong baselines. This demonstrates that the representations learned by WiFo-M2 successfully generalize to the unseen S1 scenario.
@@ V-E2 C1 Dataset
Table V summarizes the cross-scenario generalization performance of WiFo-M2 variants when applied to CE, CI, and CP modules on the unseen C1 dataset. As can be seen, the WiFo-M2 model with multi-modality provides consistent and often the best performance gains across most schemes, achieving the lowest NMSE for both CE schemes and one CP scheme (WiFo). Importantly, the WiFo-M2-Img consistently outperforms its ImageNet-pre-trained counterpart, confirming the effectiveness of the proposed ContraSoM strategy. BP results exhibit the same trend. Using only LiDAR, the MM‑BP-LiDAR baseline achieves top‑1/3/5 accuracy of $69.4\%$, $86.0\%$, and $86.0\%$, while WiFo-M2-LiDAR attains $73.1\%$, $90.3\%$, and $99.5\%$. With image alone, the vision-BP baseline delivers $78.5\%$, $88.2\%$, and $90.3\%$, against $77.4\%$, $85.5\%$, and $89.2\%$ for WiFo-M2-Img. Under multi-modal fusion, MM-BP scheme peaks at $88.7\%$, $99.5\%$, and $100\%$, whereas WiFo-M2 achieves $83.9\%$, $91.4\%$, and $93.0\%$.
TABLE V: Evaluation of the enhancement effects of WiFo-M2 models on various transceiver design schemes in C1 dataset. Values in bold indicate the best performance.
Schemes Transceiver Modules CE: CENN CE: FCDAMP CI: LPCCNet CI: SR-CI CP: WiFo CP: Transformer
@@
TABLE VI: Evaluation of the enhancement effects of WiFo-M2-Img model on various transceiver design schemes in V1 dataset. Values in bold indicate the best performance.
Schemes Transceiver Modules CE: CENN CE: FCDAMP CI: LPCCNet CI: SR-CI CP: WiFo CP: Transformer

### v2

@@ V-C1 PHY Action 1-Beam Prediction
  • Experiment Settings: The training hyperparameters for Vision-BP and MM-BP schemes are kept identical to those used in [10] and [28]. The proposed WiFo-M2-based BP scheme is trained with a batch size of $32$ and an initial learning rate of $2\times 10^{-4}$.
  • WiFo-M2’s Performance: We evaluate the performance of BP using the top-$k$ accuracy metric. This metric measures the percentage of test samples for which the ground-truth optimal beam is included in the model’s $k$ highest-ranked predictions. In our experiments, we report results for $k=1,3,5$ to comprehensively assess the prediction capability.
    As shown in Fig. 7, across all $7$ datasets, WiFo-M2 models exhibit highly competitive prediction accuracy even though the backbones are frozen. The WiFo-M2-Img scheme attains noticeably higher Top-1 prediction accuracy than Vision-BP scheme on L4, L5 and L6, whereas on the remaining four datasets its performance is only slightly lower than that of Vision-BP. This demonstrates that the WiFo-M2-Img backbone acquires channel-aware knowledge extracted during pre-training. The effect is more pronounced on LiDAR-based BP: WiFo-M2-LiDAR substantially surpasses the fully-trained MM-BP-LiDAR baseline on every dataset, which indicates that the pre-trained WiFo-M2-LiDAR captures spatial patterns that are highly predictive of the optimal beam. When both modalities are available, WiFo-M2 delivers virtually the same performance as the MM-BP scheme, with a mean Top-1 gap of only 2%. In summary, the BP results confirm that the ContraSoM pre-training aligns visual and LiDAR features with the wireless channel domain. Consequently, high-quality beam selection can be achieved with minimal retraining overhead.
V-C2 PHY Action 2-Channel Estimation
  • Problem Description: CE action refers to the process in which the BS estimates the channel $\mathbf{h}$ at the pilot-placed subcarriers. We assume that the BS adopts the hybrid (analog/digital) structure and is equipped with $N_{\rm{RF}}$ RF chains. The digital and analog precoding matrices are represented by $\mathbf{F}_{\rm{B}}\in\mathbb{C}^{N_{\rm{RF}}\times N_{\rm{RF}}}$ and $\mathbf{F}_{\rm{R}}\in\mathbb{C}^{N\times N_{\rm{RF}}}$ with $\|\mathbf{F}_{\rm{R}}\mathbf{F}_{\rm{B}}\|_{\rm{F}}^{2}=N_{\rm{RF}}$. $\mathbf{F}_{\rm{B}}$ is set as an identity matrix $\mathbf{I}_{N_{\rm{RF}}}$ during CE. The quantization bit number of the phase shifters adopted at BS is set to $B=3$. Each element of $\mathbf{F}_{\rm{R}}$ is randomly chosen from the set $\mathcal{A}_{Q}=\bigl\{\,e^{j\frac{2\pi q}{Q}}\,\bigm|\,q=1,\dots,Q\bigr\}$, where $Q\triangleq 2^{B}=8$. During CE, the single-antenna user transmits known pilots $\mathbf{x}\in\mathbb{C}^{1\times T}$ over $T$ successive time slots. The received signal sequence at BS $\mathbf{Y}\in\mathbb{C}^{N_{\rm RF}\times T}$ can then be expressed as:
@@ V-C2 PHY Action 2-Channel Estimation
    where $\mathbf{N}\in\mathbb{C}^{N\times T}$ denotes the additive white Gaussian noise with i.i.d. entries of zero mean and variance $\sigma^{2}$. Upon receiving the pilots, BS employs a certain CE network to estimate the CSI: $\hat{\mathbf{h}}=\mathcal{Q}_{\text{CE}}\bigl(\mathbf{Y},\mathcal{G}(\mathbf{M}))$, where $\mathcal{Q}_{\rm CE}(\cdot)$ represents a certain CE scheme and $\mathcal{G}(\cdot)$ denotes the WiFo‑M2 model.
@@
  • WiFo-M2’s Performance: The normalized mean square error (NMSE) is used as the estimation performance metric given by ${\rm NMSE}={{\|\hat{\mathbf{H}}-\mathbf{H}\|}_{\rm{F}}^{2}}/{{\|\mathbf{H}\|}_{\rm{F}}^{2}}$. We evaluate the channel estimation performance of CENN and FCDAMP schemes under varying SNR conditions. As shown in Fig. 8, the WiFo-M2-enhanced variants consistently outperform the original CENN and FCDAMP models across most datasets and SNRs, demonstrating the general utility of the OOB channel-aware features. Furthermore, multi-modal fusion (WiFo-M2) typically yields the most robust and significant gains, with LiDAR-only and image-only enhancements providing comparable improvements in most scenarios. On the L3 dataset with a large $256$-antenna array and limited pilots, the performance gains are less pronounced and sometimes irregular. This aligns with the expectation that under extremely sparse measurements and high-resolution array requirements, the coarse prior provided by multi-modal sensing becomes less effective in refining the estimation.
@@
V-C3 PHY Action 3-Channel Interpolation
  • Problem Description: During the CE stage, pilots are placed only on a subset of subcarriers, and a subset of BS’s antenna elements is activated [^1] In CI action, we assume that a subset of antennas are provided with coarse channel estimates, while the CI action are required to recover the full CSI across all antennas. Although this setup does not reflect a practical system configuration, it serves as an effective testbed for evaluating the evaluated CI schemes’ ability to perform spatial super‑resolution reconstruction.. Upon receiving pilots, BS estimates the channel at pilot positions using the Least Squares (LS) algorithm. All the available channel estimates form the low-resolution (LR) channel observation, denoted by $\hat{\mathbf{H}}^{\rm LR}\in\mathbb{C}^{K_{\rm p}\times N_{\rm p}}$. Specifically, if pilots are placed every $k$ subcarrier and every $x$ antenna element, the LR observation can be expressed as:
@@ V-C3 PHY Action 3-Channel Interpolation
  • Experiment Settings: We assume that the pilot placement interval in the subcarrier dimension and the sampling ratios for antenna dimensions are set to $k=K/8$ and $x=N/8$, respectively. That is, coarse channel estimates are available on only $1/8$ of the subcarriers and $1/8$ of the antennas, respectively. For network hyper-parameters, the SR-CI scheme uses a learning rate of $1\times 10^{-2}$, the LPCCNet scheme uses $1\times 10^{-4}$. The batch size for both schemes is $32$.
  • WiFo-M2’s Performance: The performance gains of WiFo-M2 models over LPCCNet and SR-CI schemes are illustrated in Fig. 9. Among the three variants, WiFo-M2 consistently achieves the most stable and significant improvements across different datasets. This demonstrates the effectiveness of multi-modal fusion in enhancing CI accuracy. Note that on the L3 dataset the WiFo-M2 models introduce a marginal performance degradation. A likely reason is that the L3 dataset configures a large antenna array, which translates into a very fine angular-resolution requirement for CI. In such a regime, the spatial side information from multi-modal sensing remain too coarse or imperfectly aligned with the complex channel, so the supplementary features cannot improve the high-precision estimates demanded by LPCCNet. Moreover, the performance gains brought by WiFo-M2 are especially pronounced on the L4 and L6 datasets. L4 employs a BS with a small antenna array, while L6 is configured with relatively fewer subcarriers. Under these reduced spatial- or frequency-resolution settings, the OOB features provided by WiFo-M2 are just sufficient to recover the correspondingly lower-dimensional CSI, leading to the sharp performance boost.
V-C4 PHY Action 4-Channel Prediction
  • Problem Description: In frequency division duplex (FDD) systems, where uplink-downlink reciprocity does not hold, the BS relies on uplink channel feedback to obtain downlink CSI. However, the feedback overhead scales linearly with the number of antennas and devices, making it infeasible for massive MIMO systems. Given that uplink and downlink transmissions experience the same physical environment, it is reasonable to utilize the uplink CSI measured at the BS to predict the downlink CSI. Similarly to [36, 37], the first half of the subcarriers is considered as the uplink channel, while the remaining half is treated as the downlink channel. Let $K_{\rm in}$ be the number of subcarriers in the first half and $K_{\rm out}$ represent the number of subcarriers in the second half. Then, the uplink channel $\mathbf{H}_{\rm d}$ and downlink channel $\mathbf{H}_{\rm u}$ can be expressed as:
@@ V-C4 PHY Action 4-Channel Prediction
    The CP action can be formulated as: $\mathbf{H}_{\rm u}=\mathcal{Q}_{\rm CP}({\mathbf{H}}_{\rm d},\mathcal{G}(\mathbf{M}))$, where $\mathcal{Q}_{\rm CP}(\cdot)$ represents a certain CP scheme.
  • Existing CP schemes: We selected two CP schemes, namely WiFo [23] and Transformer-based scheme [38]. WiFo [23] is a wireless FM designed for time/frequency-domain CP, which is pretrained via several self-supervised training tasks. The transformer-based scheme [38] is a parallel CP scheme.
@@
  • WiFo-M2’s Performance: The performance enhancements provided by WiFo-M2 models on top of baseline WiFo and Transformer-based schemes are summarized in Table III. The NMSE is also used as performance metric. The WiFo-M2 model leveraging multi-modality demonstrates its superiority by delivering highest gains for both CP baselines across most datasets, underscoring the benefit of cross-modal feature fusion. Consistent with observations in CI action, performance gains are particularly significant on datasets L4 and L6. Notably, WiFo-M2’s effectiveness is inherently bounded by the convergence of the baseline scheme itself. For instance, on dataset L5, the Transformer-based baseline fails to achieve downlink CSI prediction, and correspondingly, WiFo-M2 provides no enhancement. This indicates that the supplementary features cannot compensate for fundamental algorithmic limitations in the baseline model.
@@ V-D Ablation Study
Due to page limits, we present ablation results for several representative PHY actions, as shown in Table IV. First, removing the temporal feature extrapolation function in WiFo-M2 leads to a performance drop, confirming that extrapolating fine-grained future features effectively bridges the sparsity of sensory data. The NMSE results reported are averaged over the L1–L7 datasets. The ablation study results demonstrate the effectiveness of each proposed component. Second, replacing the ContraSoM pretrained weights with generic ImageNet‑pretrained weights significantly degrades performance (e.g., from 9.63-9.63 to 7.55-7.55 on LPCCNet), highlighting the advantage of the supplementary channel features over general‑purpose visual features. Third, removing the DA design from WiFo-M2-LiDAR causes a performance decline, validating that the proposed feature‑space diffusion process enhances LiDAR representation learning.
@@ V-E1 S1 Dataset
The S1 dataset provides measured image and LiDAR point cloud data, as well as the receive power with beams. Due to the absence of CSI, only the performance of BP action is evaluated on S1 dataset. Note that BP action on S1 dataset is relatively straightforward due to the small 16-element antenna array. The Vision-BP baseline achieves Top-1/3/5 accuracies of 91.0%, 98.7%, and 100%, while the MM-BP scheme reaches 92.2%, 100%, and 100%. Notably, the WiFo-M2-img model, with its pre-trained backbone frozen and only a final MLP layer fine-tuned, attains 91.0%, 100%, and 100%, matching or surpassing the strong baselines. This demonstrates that the representations learned by WiFo-M2 successfully generalize to the unseen S1 scenario.
@@ V-E2 C1 Dataset
Table V summarizes the cross-scenario generalization performance of WiFo-M2 variants when applied to CE, CI, and CP actions on the unseen C1 dataset. As can be seen, the WiFo-M2 model with multi-modality provides consistent and often the best performance gains across most schemes, achieving the lowest NMSE for both CE schemes and one CP scheme (WiFo). Importantly, the WiFo-M2-Img consistently outperforms its ImageNet-pre-trained counterpart, confirming the effectiveness of the proposed ContraSoM strategy. BP results exhibit the same trend. Using only LiDAR, the MM‑BP-LiDAR baseline achieves top‑1/3/5 accuracy of $69.4\%$, $86.0\%$, and $86.0\%$, while WiFo-M2-LiDAR attains $73.1\%$, $90.3\%$, and $99.5\%$. With image alone, the vision-BP baseline delivers $78.5\%$, $88.2\%$, and $90.3\%$, against $77.4\%$, $85.5\%$, and $89.2\%$ for WiFo-M2-Img. Under multi-modal fusion, MM-BP scheme peaks at $88.7\%$, $99.5\%$, and $100\%$, whereas WiFo-M2 achieves $83.9\%$, $91.4\%$, and $93.0\%$.
TABLE V: Evaluation of the enhancement effects of WiFo-M2 models on various PHY actions on C1 dataset. Values in bold indicate the best performance.
Schemes PHY Actions CE: CENN CE: FCDAMP CI: LPCCNet CI: SR-CI CP: WiFo CP: Transformer
@@
TABLE VI: Evaluation of the enhancement effects of WiFo-M2-Img model on various PHY actions on V1 dataset. Values in bold indicate the best performance.
Schemes PHY Actions CE: CENN CE: FCDAMP CI: LPCCNet CI: SR-CI CP: WiFo CP: Transformer